Eddie borrows R12000 from Janet at the beginning of 2005, at a rate of 10% p.a. compound interest. At the end of each year, he pays back R500. How much must he still pay at the end of 2008? How much does he have to pay back altogether?





1
Expert's answer
2022-03-09T10:45:42-0500

Solution


At the end of the year 2005, the payable amount is


= Principal amount + Interest amount - 500

"= 12000 + 10\% of 12000 - 500 \\

= 12000 + 1200 - 500 \\

= 12700\\"


At the end of the year 2006, the payable amount is


= payable at the end of year 2005 + Interest amount - 500

"= 12700 + 10\% of 12700 - 500 \\

= 12700 + 1270 - 500 \\

= 13470"


At the end of the year 2007, the payable amount is


= payable at the end of year 2006 + Interest amount - 500

"= 13470 + 10\% of 13470 - 500 \\

= 13470 + 1347 - 500 \\

= 14317"


At the end of the year 2008, the payable amount is


= payable at the end of year 2007 + Interest amount - 500

"= 14317 + 10\% of 14317 - 500 \\

= 14317 + 1431.7 - 500 \\

= 15248.7"


Hence the payable amount at the end of the year 2008, is "R = 15248.7"